Home‎ > ‎Crew 164‎ > ‎

0224-DailySummaryReport

DAILY SUMMARY REPORT
SOL: 3
Person filling out Report: Mohammad IRANMANESH, XO
Summary Title: Crazy how much work you can do in a Martian day.
Mission Status: All systems GO except ATVs. All experiments are
successfully ongoing except MOMA.
Sol Activity Summary: Morning sport routine. EVA #3. Work on projects.
EVA#3b to test ATVs with new instructions and get the supplies from
the Phobos Shuttle. Opening 'geology' experiment and reading provided
materials.
Look Ahead Plan: First night at the observatory for calibration if
good weather. With the tomorrow's EVA, most of our technology projects
would have proven their feasibility in martian conditions and we will
be able to continue with later phases of the experiments. A safety and
security meeting with the crew is schedule tonight.
Anomalies in work: Some ATVs still not working (see Engineering report).
Weather: sunny, low wind. Very hot in spacesuit about the end of morning.
Crew Physical Status: Everyone feels great.
EVA: EVA to explore south of the HAB. Test #2 of CRV and Test #3 of EMUI.
Reports to be filed:
- Commander report
- Engineering report
- Journalist report
- Health and Safety report
- EVA #4 request
- EVA #3 report
Support Requested:
- Fix remaining ATVs
- New radios